jayley14:

A bike wheel is 26 inches in diameter. What is the bike wheel's diameter in millimeters (1 inch = 25.4 millimeters)?

Shadow:

\[\frac{ 26 inches }{ 1 } \times \frac{ 25.4 millimeters }{ 1 inch}\] Multiply across, inches cancel out, millimeters stays, just compute the product of the coefficients (26 times 25.4)